OpenStudy (anonymous):

What is the value of x? Round your answer to the nearest tenth if necessary. Right triangle whose two legs measure fourteen and x, and whose hypotenuse measures twenty

OpenStudy (anonymous):



OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

20^2=14^2+x^2 x^2=204 now sqrt204 and get the answer

OpenStudy (anonymous):

x=14.282857

OpenStudy (anonymous):

14.2?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

no 14.3 as 8 > 5

OpenStudy (anonymous):

ok i see thanks

OpenStudy (anonymous):

q